# Variations of metric that preserve a Riemannian submersion and geometry of its fibers

### v1

@@ 3.1 Preserving Riemannian submersions
We note that Riemannian metrics obtained as solutions of systems (4)-(6) and (9)-(9) do not depend on the choice of adapted $g_{0}$-orthonormal frame $\{E_{a},{\cal E}_{i}\}$. Indeed, any two such frames $\{E_{a},{\cal E}_{i}\}$ and $\{E^{\prime}_{c},{\cal E}^{\prime}_{k}\}$ are related by a $g_{0}$-orthogonal transformation preserving each of distributions $\ker\pi_{*}$ and ${\cal H}(0)$, so ${\cal E}^{\prime}_{k}=\sum\nolimits_{i=n+1}^{n+p}M_{ik}{\cal E}_{i}$, $E^{\prime}_{c}=\sum\nolimits_{a=1}^{n}M_{ac}E_{a}$, with $M_{\mu\nu}$ being coefficients of a matrix $M\in O(n)\times O(p)$. Hence, e.g., $g^{\prime}_{kc}=\sum\nolimits_{i=n+1}^{n+p}\sum\nolimits_{a=1}^{n}M_{ik}M_{ac}g_{ai}$ and, by (11), we have $\lambda^{\prime}_{ck}=\sum\nolimits_{i=n+1}^{n+p}\sum\nolimits_{a=1}^{n}\lambda_{ai}M_{ik}M_{ac}$. Comparing systems (4)-(6) (and similarly, (9)-(9)) written with and without primed coefficients, we get $\partial_{t}g^{\prime}_{kc}=\sum\nolimits_{i=n+1}^{n+p}\sum\nolimits_{a=1}^{n}\partial_{t}(M_{ik}M_{ac}g_{ai})$ and $\partial_{t}g^{\prime}_{kl}=\sum\nolimits_{i=n+1}^{n+p}\sum\nolimits_{j=n+1}^{n+p}\partial_{t}(M_{ik}M_{jl}g_{ij})$, so from the uniqueness of solution of ODE system, $g^{\prime}_{kc}(t)=\sum\nolimits_{i=n+1}^{n+p}\sum\nolimits_{a=1}^{n}M_{ik}M_{ac}g_{ai}(t)$ and $g^{\prime}_{kl}(t)=\sum\nolimits_{i=n+1}^{n+p}\sum\nolimits_{j=n+1}^{n+p}M_{ik}M_{jl}g_{ij}(t)$ for all $t$ for which the solution exist.
Corollary 3.6.
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion. For every pair of sets: $\{V_{n+1},\ldots,V_{n+p}\}$ of vertical vector fields, bounded on $(M,g_{0})$, and $\{W_{n+1},\ldots,W_{n+p}\}$ of linearly independent vector fields, bounded on $(N,g_{N})$, there exists a unique variation $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,\ker\pi_{*},g_{0})$ that preserves the Riemannian submersion $\pi$, such that for every $X\in\mathfrak{X}_{\ker\pi_{*}}$
@@
Let $\{E_{a},{\cal E}_{i}\}$ be a local adapted $g_{0}$-orthonormal frame on an open set $U\subset M$. Then there exist functions $\{\lambda_{ai}:U\times\mathbb{R}\rightarrow\mathbb{R}\}$ such that
@@
For these $\{\lambda_{ai}\}$, let $g_{t}$ be the solution of (9)-(9). Then, from (9) it follows that coefficients $\{b_{ai}\}$ of $B_{t}=\partial_{t}g_{t}$ are given by (10). From (9) we obtain that $B_{t}^{\sharp}E_{a}$ are $g_{t}$-horizontal, hence
$B^{\sharp}_{t}E_{a}=\sum\nolimits_{i=n+1}^{n+p}\left({\tilde{\lambda}}_{ai}{\cal E}_{i}-\sum\nolimits_{c=1}^{n}g_{ic}{\tilde{\lambda}}_{ai}E_{c}\right)$ (14)
@@
for some $\{\tilde{\lambda}_{ai}\}$. But obtaining $b_{ai}=B_{t}(E_{a},{\cal E}_{i})=g_{t}(B^{\sharp}_{t}E_{a},{\cal E}_{i})$ from (14) and comparing with (10), we get for all $a\in\{1,\ldots,n\}$ and all $i\in\{n+1,\ldots,n+p\}$
$\sum\nolimits_{j=n+1}^{n+p}({\tilde{\lambda}}_{aj}-\lambda_{aj})(g_{ij}-\sum\nolimits_{b=1}^{n}g_{jb}g_{bi})=0.$ (15)
@@
Since $g_{ij}-\sum\nolimits_{b=1}^{n}g_{jb}g_{bi}=g_{t}(P_{{\cal H}}^{t}{\cal E}_{i},P_{{\cal H}}^{t}{\cal E}_{j})$ and $\{P_{{\cal H}}^{0}{\cal E}_{i}\}$ are linearly independent, there exists $\epsilon^{\prime}>0$ such that for $t\in(-\epsilon^{\prime},\epsilon^{\prime})$ we have $\det g_{t}(P_{{\cal H}}^{t}{\cal E}_{i},P_{{\cal H}}^{t}{\cal E}_{j})\neq 0$ and from (15) it follows ${\tilde{\lambda}}_{ai}=\lambda_{ai}$. Hence, (11) holds and, by (13), yields (12).
@@
Now we prove that there exists $\epsilon>0$ such that at every $x\in M$ the solution of (9)-(9) exists for all $t\in(-\epsilon,\epsilon)$. Indeed, (9)-(9) are equations for coefficients of $g_{t}$ in a $g_{0}$-orthonormal frame, so the initial condition for these equations for all $x\in M$ is the same: $g_{ai}(0)=0$ and $g_{ij}(0)=\delta_{ij}$ for $a\in\{1,\ldots,n\}$ and $i,j\in\{n+1,\ldots,n+p\}$. Since $\{V_{i},W_{i}\}$ are bounded, from (13) it follows that there exists a compact set $K\subset\mathbb{R}^{np}$, such that $\{\lambda_{ai}\}\in K$ for all $x\in M$, and for all local $g_{0}$-orthonormal frames $\{E_{a},{\cal E}_{i}\}$. For every $\{\lambda_{ai}\}\in K$ there exist $\epsilon^{\prime}>0$ and an open neighbourhood ${\cal N}\subset\mathbb{R}^{np}$ such that for all $\{\lambda_{ai}\}\in{\cal N}$ the solution of (9)-(9) with initial condition $g_{ai}(0)=0,g_{ij}(0)=\delta_{ij}$ exists and, since the solutions depend smoothly on parameters $\{\lambda_{ai}\}$ of the equations, $\det(g_{ij}(t)-\sum\nolimits_{b=1}^{n}g_{jb}(t)g_{bi}(t))\neq 0$ for all $t\in(-\epsilon^{\prime},\epsilon^{\prime})$. Covering $K$ with these neighbourhoods and finding the smallest $\epsilon^{\prime}$ for some finite subcover, we obtain an interval $(-\epsilon,\epsilon)$ on which $g_{t}$ is defined for all $x\in M$ as the metric whose components in a local frame $\{E_{a},{\cal E}_{i}\}$ are the solution of (9)-(9) with $\{\lambda_{ai}\}$ defined by (13).
Since $\{\lambda_{ai}\}$ are obtained from globally defined vector fields in (13), solutions obtained locally in adapted $g_{0}$-orthonormal frames give rise to a tensor field $g_{t}$ on $M$. ∎
@@
Let $U\subset N$ be an open set, on which linearly independent vector fields $W_{n+1},\ldots,W_{n+p}$ are defined. Then on $\pi^{-1}(U)$ every one-parameter family of metrics $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,\ker\pi_{*},g_{0})$, that preserves the Riemannian submersion $\pi$, satisfies (12) for some $t$-dependent vertical vector fields $\{V_{n+1}(t),\ldots,V_{n+p}(t)\}$ on $M$. This follows from $B_{t}^{\sharp}X$ being $g_{t}$-horizontal for every vertical $X$ and $\{P_{{\cal H}}^{t}\pi^{*}W_{n+1},\ldots,P_{{\cal H}}^{t}\pi^{*}W_{n+p}\}$ being a linear frame of distribution ${\cal H}(t)$ on $\pi^{-1}(U)$.
@@
[14] Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion and let $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,\ker\pi_{*},g_{0})$ be a variation. If $\{E_{a},{\cal E}_{i}\}$ is a local $g_{0}$-orthonormal frame on $M$ such that $E_{a}\in\mathfrak{X}_{\ker\pi_{*}}$ for all $a\in\{1,\ldots,n\}$, then $\{E_{a},{\cal E}_{i}(t)\}$ such that
@@
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion and let $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,\ker\pi_{*},g_{0})$ be a variation. Then for all $X,Y,Z\in\mathfrak{X}_{\ker\pi_{*}}$ we have
@@
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion and let variation $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,\ker\pi_{*},g_{0})$ preserve the Riemannian submersion $\pi$. Then for all $X,Y\in\mathfrak{X}_{\ker\pi_{*}}$ and $Z\in\mathfrak{X}_{{\cal H}(t)}$ we have
@@
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion and let variation $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,\ker\pi_{*},g_{0})$ preserve the Riemannian submersion $\pi$. Then for every local $g_{0}$-orthonormal frame $\{E_{1},\ldots,E_{n}\}\subset\mathfrak{X}_{\ker\pi_{*}}$ and for all $Z\in\mathfrak{X}_{{\cal H}(t)}$ we have
@@ 3.2 Preserving extrinsic geometry of fibers
We will write that a vector field is non-trivial if it is not everywhere zero, and that a variation $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,\ker\pi_{*},g_{0})$ is non-trivial if $g_{t}\neq g_{0}$ for some $t\in(-\epsilon,\epsilon)$.
@@
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion with totally geodesic fibers. If there exists a non-trivial variation $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,\ker\pi_{*},g_{0})$ such that $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is a Riemannian submersion with totally geodesic fibers for all $t\in(-\epsilon,\epsilon)$, then there exists a non-trivial vertical vector field $\xi$, whose restriction to $({\cal F}_{x},g_{0}|_{{\cal F}_{x}})$ is a Killing field for all $x\in N$. On the other hand, if on $(M,g_{0})$ exists a non-trivial bounded vertical vector field $\xi$, whose restriction to every fiber of $\pi$ is Killing, then there exists a non-trivial variation $g_{t}$ with the above properties.
@@
Let $x\in N$. Let $U\subset N$ be a neighbourhood of $x$ on which there exist linearly independent vector fields $\{W_{n+1},\ldots,W_{n+p}\}$, then, by Corollary 3.6 and Remark 3.7, we can assume that (12) holds on $\pi^{-1}(U)$. We will restrict our considerations to $\pi^{-1}(U)$.
For all $X\in\mathfrak{X}_{\ker\pi_{*}}$ and all $i\in\{n+1,\ldots,n+p\}$, since $\pi^{*}W_{i}$ is projectable, we have $P_{{\cal H}}^{t}[X,\pi^{*}W_{i}]=0$ and since $\ker\pi_{*}$ is integrable, we have $P_{{\cal H}}^{t}[X,P_{{\cal V}}^{t}\pi^{*}W_{i}]=0$, so
@@
Hence, from (19) and (12) we obtain
$2P_{{\cal H}}^{t}\partial_{t}\nabla^{t}_{X}Y=P_{{\cal H}}^{t}\big{(}[X,\sum\nolimits_{i=n+1}^{n+p}g_{0}(V_{i},Y)P_{{\cal H}}^{t}\pi^{*}W_{i}]$
$\quad+[Y,\sum\nolimits_{i=n+1}^{n+p}g_{0}(V_{i},X)P_{{\cal H}}^{t}\pi^{*}W_{i}]-B_{t}^{\sharp}P_{{\cal V}}^{0}(\nabla^{0}_{X}Y+\nabla^{0}_{Y}X)\big{)}$
$=\sum\nolimits_{i=n+1}^{n+p}X(g_{0}(V_{i},Y))P_{{\cal H}}^{t}\pi^{*}W_{i}+\sum\nolimits_{i=n+1}^{n+p}Y(g_{0}(V_{i},X))P_{{\cal H}}^{t}\pi^{*}W_{i}$
$\quad-\sum\nolimits_{i=n+1}^{n+p}g_{0}(V_{i},\nabla^{0}_{X}Y+\nabla^{0}_{Y}X)P_{{\cal H}}^{t}\pi^{*}W_{i}$
$=\sum\nolimits_{i=n+1}^{n+p}\big{(}g_{0}(\nabla^{0}_{X}V_{i},Y)+g_{0}(\nabla^{0}_{Y}V_{i},X)\big{)}P_{{\cal H}}^{t}\pi^{*}W_{i}.$ (25)
The fibers of $\pi:(M,g_{t})\rightarrow(N,g_{N})$ remain totally geodesic if and only if $P_{{\cal H}}^{t}\partial_{t}\nabla^{t}_{X}Y=0$.
Suppose that $g_{t}\neq g_{0}$ for some $t\in(-\epsilon,\epsilon)$ at some $y\in{\cal F}_{x}$. If all $V_{i}=0$, then $g_{t}=g_{0}$, so to have a non-trivial variation $g_{t}$ such that $P_{{\cal H}}^{t}\partial_{t}\nabla^{t}_{X}Y=0$ for all $X,Y\in\mathfrak{X}_{\ker\pi_{*}}$, at least one non-trivial vector field $\xi\in\{V_{n+1},\ldots,V_{n+p}\}$ must satisfy

### v2

@@ 3.1 Preserving Riemannian submersions
We note that Riemannian metrics obtained as solutions of system (9)-(9) do not depend on the choice of adapted $g_{0}$-orthonormal frame $\{E_{a},{\cal E}_{i}\}$ if coefficients $\{\lambda_{ai}\}$ are properly transformed when changing the frame. Indeed, any two such frames $\{E_{a},{\cal E}_{i}\}$ and $\{E^{\prime}_{c},{\cal E}^{\prime}_{k}\}$ are related by a $g_{0}$-orthogonal transformation preserving each of distributions $\cal V$} and ${\cal H}(0)$, so ${\cal E}^{\prime}_{k}=\sum\nolimits_{i=n+1}^{n+p}M_{ik}{\cal E}_{i}$, $E^{\prime}_{c}=\sum\nolimits_{a=1}^{n}M_{ac}E_{a}$, with $M_{\mu\nu}$ being coefficients of a matrix $M\in O(n)\times O(p)$. Hence, e.g., $g^{\prime}_{kc}=\sum\nolimits_{i=n+1}^{n+p}\sum\nolimits_{a=1}^{n}M_{ik}M_{ac}g_{ai}$. Let the transformation rule for $\{\lambda_{ai}\}$ be the following:
$\lambda^{\prime}_{ck}=\sum\nolimits_{i=n+1}^{n+p}\sum\nolimits_{a=1}^{n}\lambda_{ai}M_{ik}M_{ac}.$ (11)
Comparing the system (9)-(9) written in frames $\{E_{a},{\cal E}_{i}\}$ and $\{E^{\prime}_{c},{\cal E}^{\prime}_{k}\}$, we get $\partial_{t}g^{\prime}_{kc}=\sum\nolimits_{i=n+1}^{n+p}\sum\nolimits_{a=1}^{n}\partial_{t}(M_{ik}M_{ac}g_{ai})$ and $\partial_{t}g^{\prime}_{kl}=\sum\nolimits_{i=n+1}^{n+p}\sum\nolimits_{j=n+1}^{n+p}\partial_{t}(M_{ik}M_{jl}g_{ij})$, so from the uniqueness of solution of ODE system, we have $g^{\prime}_{kc}(t)=\sum\nolimits_{i=n+1}^{n+p}\sum\nolimits_{a=1}^{n}M_{ik}M_{ac}g_{ai}(t)$ and $g^{\prime}_{kl}(t)=\sum\nolimits_{i=n+1}^{n+p}\sum\nolimits_{j=n+1}^{n+p}M_{ik}M_{jl}g_{ij}(t)$ for all $t$, for which the solution exists.
Theorem 3.6.
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion. For every pair of sets: $\{V_{n+1},\ldots,V_{n+p}\}$ of vertical vector fields, bounded on $(M,g_{0})$, and $\{W_{n+1},\ldots,W_{n+p}\}$ of linearly independent vector fields, bounded on $(N,g_{N})$, there exists a unique variation $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$ that preserves the Riemannian submersion $\pi$, such that for every $X\in\mathfrak{X}_{{\cal V}}$
@@
Let $\{E_{a},{\cal E}_{i}\}$ be a local adapted $g_{0}$-orthonormal frame on an open set $U^{\prime}\subset M$, and let $U$ be a relatively compact, open set, such that $\overline{U}\subset U^{\prime}$. Then there exist functions $\{\lambda_{ai}:U^{\prime}\rightarrow\mathbb{R}\}$ such that
@@
For these $\{\lambda_{ai}\}$, let $\{g_{t},t\in(-\epsilon^{\prime},\epsilon^{\prime})\}\subset\operatorname{Riem}(U,{\cal V},g_{0})$ be the variation obtained from Lemma 3.4.
@@
From (9) it follows that coefficients $\{b_{ai}\}$ of $B_{t}=\partial_{t}g_{t}$ are given by (10). From (9) we obtain that $B_{t}^{\sharp}E_{a}$ are $g_{t}$-horizontal, hence
$B^{\sharp}_{t}E_{a}=\sum\nolimits_{i=n+1}^{n+p}\left({\tilde{\lambda}}_{ai}(t){\cal E}_{i}-\sum\nolimits_{c=1}^{n}g_{ic}(t){\tilde{\lambda}}_{ai}(t)E_{c}\right)$ (14)
@@
for some $\{\tilde{\lambda}_{ai}:U\times\mathbb{R}\rightarrow\mathbb{R}\}$. But obtaining $b_{ai}=B_{t}(E_{a},{\cal E}_{i})=g_{t}(B^{\sharp}_{t}E_{a},{\cal E}_{i})$ from (14) and comparing with (10), we get for all $a\in\{1,\ldots,n\}$ and all $i\in\{n+1,\ldots,n+p\}$
$\sum\nolimits_{j=n+1}^{n+p}({\tilde{\lambda}}_{aj}(t)-\lambda_{aj})(g_{ij}(t)-\sum\nolimits_{b=1}^{n}g_{jb}(t)g_{bi}(t))=0.$ (15)
@@
Since $g_{ij}(t)-\sum\nolimits_{b=1}^{n}g_{jb}(t)g_{bi}(t)=g_{t}(P_{{\cal H}}^{t}{\cal E}_{i},P_{{\cal H}}^{t}{\cal E}_{j})$ and, by Lemma 3.4, $\{P_{{\cal H}}^{t}{\cal E}_{i}\}$ are linearly independent, for all $t\in(-\epsilon^{\prime},\epsilon^{\prime})$ we have $\det g_{t}(P_{{\cal H}}^{t}{\cal E}_{i},P_{{\cal H}}^{t}{\cal E}_{j})\neq 0$ and from (15) it follows that ${\tilde{\lambda}}_{ai}=\lambda_{ai}$. Hence,
$B^{\sharp}_{t}E_{a}=\sum\nolimits_{i=n+1}^{n+p}\lambda_{ai}\left({\cal E}_{i}-\sum\nolimits_{c=1}^{n}g_{ic}(t)E_{c}\right)$ (16)
holds and, by (13), yields (12).
@@
Now we prove that there exists $\epsilon>0$ such that at every $x\in M$ the solution of (9)-(9) exists for all $t\in(-\epsilon,\epsilon)$. Indeed, (9)-(9) are equations for coefficients of $g_{t}$ in a $g_{0}$-orthonormal frame, so the initial condition for these equations for all $x\in M$ is the same: $g_{ai}(0)=0$ and $g_{ij}(0)=\delta_{ij}$ for $a\in\{1,\ldots,n\}$ and $i,j\in\{n+1,\ldots,n+p\}$. Since $\{V_{i},W_{i}\}$ are bounded, from (13) it follows that there exists a compact set $K\subset\mathbb{R}^{np}$, such that $\{\lambda_{ai}\}\in K$ for all $x\in M$, and for all local $g_{0}$-orthonormal frames $\{E_{a},{\cal E}_{i}\}$. For every $\{\lambda_{ai}\}\in K$ there exist $\epsilon^{\prime}>0$ and an open neighbourhood ${\cal N}\subset\mathbb{R}^{np}$ such that for all $\{\lambda_{ai}\}\in{\cal N}$ the solution of (9)-(9) with initial condition $g_{ai}(0)=0,g_{ij}(0)=\delta_{ij}$ exists and, since the solutions depend smoothly on parameters $\{\lambda_{ai}\}$ of the equations, $\det(g_{ij}(t)-\sum\nolimits_{b=1}^{n}g_{jb}(t)g_{bi}(t))\neq 0$ for all $t\in(-\epsilon^{\prime},\epsilon^{\prime})$. Covering $K$ with these neighbourhoods and finding the smallest $\epsilon^{\prime}$ for some finite subcover, we obtain an interval $(-\epsilon,\epsilon)$ on which $g_{t}$ is defined for all $x\in M$ as the metric whose components in a local frame $\{E_{a},{\cal E}_{i}\}$ are the solution of (9)-(9) with $\{\lambda_{ai}\}$ defined by (13). By Remark 3.5, since $\{\lambda_{ai}\}$ defined by (13) satisfy the transformation rule (11) and are obtained from globally defined vector fields, solutions obtained locally in adapted $g_{0}$-orthonormal frames give rise to a tensor field $g_{t}$ on $M$. ∎
@@
Let $U\subset N$ be an open set, on which linearly independent vector fields $W_{n+1},\ldots,W_{n+p}$ are defined. Then on $\pi^{-1}(U)$ every one-parameter family of metrics $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$, that preserves the Riemannian submersion $\pi$, satisfies (12) for some $t$-dependent vertical vector fields $\{V_{n+1}(t),\ldots,V_{n+p}(t)\}$ on $M$. This follows from $B_{t}^{\sharp}X$ being $g_{t}$-horizontal for every vertical $X$ and $\{P_{{\cal H}}^{t}\pi^{*}W_{n+1},\ldots,P_{{\cal H}}^{t}\pi^{*}W_{n+p}\}$ being a linear frame of distribution ${\cal H}(t)$ on $\pi^{-1}(U)$.
@@
[16] Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion and let $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$ be a variation. If $\{E_{a},{\cal E}_{i}\}$ is a local $g_{0}$-orthonormal frame on $M$ such that $E_{a}\in\mathfrak{X}_{{\cal V}}$ for all $a\in\{1,\ldots,n\}$, then $\{E_{a}(t),{\cal E}_{i}(t)\}$ such that
@@
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion and let $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$ be a variation. Then for all $X,Y,Z\in\mathfrak{X}_{{\cal V}}$ we have
@@
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion and let variation $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$ preserve the Riemannian submersion $\pi$. Then for all $X,Y\in\mathfrak{X}_{{\cal V}}$ and $Z\in\mathfrak{X}_{{\cal H}(t)}$ we have
@@
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion and let variation $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$ preserve the Riemannian submersion $\pi$. Then for every local $g_{0}$-orthonormal frame $\{E_{1},\ldots,E_{n}\}\subset\mathfrak{X}_{{\cal V}}$ and for all $Z\in\mathfrak{X}_{{\cal H}(t)}$ we have
@@ 3.2 Preserving extrinsic geometry of fibers
We will write that a vector field is non-trivial if it is not everywhere zero, and that a variation $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$ is non-trivial if $g_{t}\neq g_{0}$ for some $t\in(-\epsilon,\epsilon)$.
@@
Let $\pi:(M,g_{0})\rightarrow(N,g_{N})$ be a Riemannian submersion with totally geodesic fibers. If there exists a non-trivial variation $\{g_{t},t\in(-\epsilon,\epsilon)\}\subset\operatorname{Riem}(M,{\cal V},g_{0})$ such that $\pi:(M,g_{t})\rightarrow(N,g_{N})$ is a Riemannian submersion with totally geodesic fibers for all $t\in(-\epsilon,\epsilon)$, then there exists a non-trivial vertical vector field $\xi$, whose restriction to $({\cal F}_{x},g_{0}|_{{\cal F}_{x}})$ is a Killing field for all $x\in N$. On the other hand, if on $(M,g_{0})$ exists a non-trivial bounded vertical vector field $\xi$, whose restriction to every fiber of $\pi$ is Killing, then there exists a non-trivial variation $g_{t}$ with the above properties.
@@
Let $x\in N$. Let $U\subset N$ be a neighbourhood of $x$ on which there exist linearly independent vector fields $\{W_{n+1},\ldots,W_{n+p}\}$, then, by Remark 3.7, we can assume that (12) holds on $\pi^{-1}(U)$. We will restrict our considerations to $\pi^{-1}(U)$.
For all $X\in\mathfrak{X}_{{\cal V}}$ and all $i\in\{n+1,\ldots,n+p\}$, since $\pi^{*}W_{i}$ is projectable, we have $P_{{\cal H}}^{t}[X,\pi^{*}W_{i}]=0$ and since $\cal V$} is integrable, we have $P_{{\cal H}}^{t}[X,P_{{\cal V}}^{t}\pi^{*}W_{i}]=0$, so
@@
Hence, from (20) and (12) we obtain for all $X,Y\in\mathfrak{X}_{{\cal V}}$
$2P_{{\cal H}}^{t}\partial_{t}\nabla^{t}_{X}Y=P_{{\cal H}}^{t}\big{(}[X,\sum\nolimits_{i=n+1}^{n+p}g_{0}(V_{i}(t),Y)P_{{\cal H}}^{t}\pi^{*}W_{i}]$
$\quad+[Y,\sum\nolimits_{i=n+1}^{n+p}g_{0}(V_{i}(t),X)P_{{\cal H}}^{t}\pi^{*}W_{i}]-B_{t}^{\sharp}P_{{\cal V}}^{0}(\nabla^{0}_{X}Y+\nabla^{0}_{Y}X)\big{)}$
$=\sum\nolimits_{i=n+1}^{n+p}X(g_{0}(V_{i}(t),Y))P_{{\cal H}}^{t}\pi^{*}W_{i}+\sum\nolimits_{i=n+1}^{n+p}Y(g_{0}(V_{i}(t),X))P_{{\cal H}}^{t}\pi^{*}W_{i}$
$\quad-\sum\nolimits_{i=n+1}^{n+p}g_{0}(V_{i}(t),\nabla^{0}_{X}Y+\nabla^{0}_{Y}X)P_{{\cal H}}^{t}\pi^{*}W_{i}$
$=\sum\nolimits_{i=n+1}^{n+p}\big{(}g_{0}(\nabla^{0}_{X}V_{i}(t),Y)+g_{0}(\nabla^{0}_{Y}V_{i}(t),X)\big{)}P_{{\cal H}}^{t}\pi^{*}W_{i}.$ (26)
The fibers of $\pi:(M,g_{t})\rightarrow(N,g_{N})$ remain totally geodesic if and only if $P_{{\cal H}}^{t}\partial_{t}\nabla^{t}_{X}Y=0$ for all $X,Y\in\mathfrak{X}_{{\cal V}}$. Indeed, using a frame from Lemma 3.8 we obtain
$\partial_{t}P_{{\cal H}}^{t}\nabla^{t}_{X}Y=P_{{\cal H}}^{t}\partial_{t}\nabla^{t}_{X}Y-\sum\nolimits_{a=1}^{n}\sum\nolimits_{i=n+1}^{n+p}g_{t}(\nabla^{t}_{X}Y,{\cal E}_{i})B_{t}({\cal E}_{i},E_{a})E_{a}.$
On the other hand, using (17) and (1), we obtain
$P_{{\cal V}}^{t}\partial_{t}\nabla^{t}_{X}Y$ $=$ $\sum\nolimits_{a=1}^{n}g_{t}(\partial_{t}\nabla^{t}_{X}Y,E_{a})E_{a}=-\sum\nolimits_{a=1}^{n}B_{t}(\nabla^{t}_{X}Y,E_{a})E_{a}$
$=$ $-\sum\nolimits_{a=1}^{n}\sum\nolimits_{i=n+1}^{n+p}g_{t}(\nabla^{t}_{X}Y,{\cal E}_{i})B_{t}({\cal E}_{i},E_{a})E_{a}.$
It follows that $\partial_{t}P_{{\cal H}}^{t}\nabla^{t}_{X}Y=P_{{\cal H}}^{t}\partial_{t}\nabla^{t}_{X}Y+P_{{\cal V}}^{t}\partial_{t}\nabla^{t}_{X}Y$. Hence, if $\partial_{t}h_{t}(X,Y)=0$ then $P_{{\cal H}}^{t}\partial_{t}\nabla^{t}_{X}Y=0$. On the other hand, by (1) and (18)
$\partial_{t}g_{t}(h_{t}(X,Y),h_{t}(X,Y))$ $=$ $B_{t}(h_{t}(X,Y),h_{t}(X,Y))+2g_{t}(\partial_{t}P_{{\cal H}}^{t}\nabla^{t}_{X}Y,P_{{\cal H}}^{t}\nabla^{t}_{X}Y)$
$=$ $2g_{t}(P_{{\cal H}}^{t}\partial_{t}\nabla^{t}_{X}Y,P_{{\cal H}}^{t}\nabla^{t}_{X}Y)-2g_{t}(P_{{\cal H}}^{t}\partial_{t}P_{{\cal V}}^{t}\nabla^{t}_{X}Y,P_{{\cal H}}^{t}\nabla^{t}_{X}Y)$
$=$ $2g_{t}(P_{{\cal H}}^{t}\partial_{t}\nabla^{t}_{X}Y,P_{{\cal H}}^{t}\nabla^{t}_{X}Y).$
Hence, if $P_{{\cal H}}^{t}\partial_{t}\nabla^{t}_{X}Y=0$ and $h_{0}(X,Y)=0$, then $h_{t}(X,Y)=0$.
Suppose that $g_{t}\neq g_{0}$ for some $t\in(-\epsilon,\epsilon)$ at some $y\in{\cal F}_{x}$. If all $V_{i}(t)=0$ for all $t\in(-\epsilon,\epsilon)$, then $g_{t}=g_{0}$, so to have a non-trivial variation $g_{t}$ such that $P_{{\cal H}}^{t}\partial_{t}\nabla^{t}_{X}Y=0$ for all $X,Y\in\mathfrak{X}_{{\cal V}}$, at least one non-trivial vector field $\xi\in\{V_{n+1}(t),\ldots,V_{n+p}(t),t\in(-\epsilon,\epsilon)\}$ must satisfy